OpenStudy (anonymous):

Use the discriminant to determine the number and type of solutions the equation has. x^2 + 8x + 10 = 0 A. no real solutions B. one real solution C. two rational solutions D. two irrational solutions

OpenStudy (anonymous):

The discriminate is the portion of the quadratic equation that is under the radical (square root). It tells you if you have real or imaginary roots because if the number under the radical is negative, you will have two imaginary roots. If the number is positive then you will have two real roots, and if the number is zero, you will have one root. (This last case means you're vertex is on the x-axis somewhere.) So in your case: a=1 b=8 c=16 Plug it into b^2 - 4ac = 64-64 = 0.

OpenStudy (anonymous):

so its B